OpenStudy (anonymous):

I need help! A ceiling fan with 81-cm diameter blades is turning @ 70rpm. Suppose the fan coasts to a stop 30s after being turned off. Through how many revolutions does the fan turn while stopping I already solved for v v=2.0m/s

OpenStudy (anonymous):

find the angular acc. then use the displacement equation

OpenStudy (anonymous):

use this to find acc (alp\[\omega _{f}=\omega _{i}+\alpha t\]ha) then use this to find theta (displacemnet) \[\theta=\omega _{i}t+\frac{ 1 }{ 2 }\alpha t ^{2}\]

OpenStudy (anonymous):

Thank you so much @AJ01!!